Home‎ > ‎Crew 149‎ > ‎

210-CommandersReport

Commander's Report
2/10/2015
Paul Bakken

Preface note:  I have been trying -- unsuccessfully -- since approximately 19:10 hours to submit this report.  The Hab Internet keeps timing out on both email submission and retrieval requests.  If I actually manage to push this report through, please be advised that I have not seen any emails directed to me during this comms window yet.

1 -- Reports incoming
2 -- Loss of power adaptations
3 -- Power restored
4 -- Kneepads received
5 -- Journalist visits
6 -- Airlock prototype system installed
7 -- Plan for Tomorrow

1. The following reports/requests will be submitted tonight, in addition to this Commander's report:

Engineering report
HSO report
Science report
GreenHab report
Journalist report

2. The first part of the day was consumed with adapting to the loss of power.  Crew spent the first few hours of the day salvaging/recovering projects that were dependent on power, water, or heat.  

Most notably, the regolith plant growth study specimens were moved to a calibrated distance away from the propane heater, which functions independent of power, on Deck 1 to maintain a temperature of no less than 70 degrees Fahrenheit.  

We also continued strict water rationing, as we did not want to draw off of the loft tank without the benefit of the pump.  We were concerned that doing so might introduce an air bubble into the line.  Tank water was reserved for use as a last resort.

We also took the step of turning off all lights and unplugging heavy use devices such as the refrigerator and computers.  We did not want utilities left on to create a sudden demand surge to the generator once power was restored to the Hab.

Finally, we made use of an AC adapter in a personally owned rover (Ken's truck) to provide power to charge personal electronics.  We performed a series of short EVAs to and from the "power rover" and the Hab to shuttle devices out to be charged and then back in again.  The primary reason for this was to keep certain cell phone charged, as they were our only viable connection with Mission Control at the time.  We also charged up a number of personally owned "charge bricks" to ensure a backup reserve of charging power, in the event that the power rover solution no longer worked.

3. Power to the Hab was restored at approximately 12:20 hours.  All subsystems were gradually and progressively brought back online, and devices such as the fridge and computers plugged back in.  The crew spent a considerable amount of time confirming the correct operation of each electrical system in the Hab after re-powering.  The electrical devices are confirmed to be working at least as well as they were before the loss of power.  We continue to struggle with slow, intermittent Internet connection, however.

4. DG Interplanetary sent a cargo drone to the Hab today containing the UND kneepads for EVA testing.  I have emailed the organization contact to let them know that we got them and will proceed with the testing. (Yes! One email actually got though!)

5. We had two visitors today, whom the crew accommodated without breaking Sim.  The Journalist report will provide more details about these visits.

6. In conjunction with the power rover EVAs, we were able to get the prototype airlock system installed.  We will send pictures soon.  For the time being, please rest assured that all components have been mounted with either Command Strips or low adhesive/non-marking tape.  We have not made any permanent alterations to the Hab.

The system works great, by the way.  The crew has remarked that it adds a lot to the veracity of Sim.

7 -- Tomorrow, Crew 149 plans:

a) We anticipate one non-engineering EVA, for which a formal request should be submitted before the end of the comms window.
b) We have activated the randomized emergency simulation application, and will continue to run it for the remainder of Sim.  
c) We plan to set up the ozone machine in the engineering airlock for testing later this week.  
d) We will also begin the emergency/transitional foods evaluation at breakfast.

The Internet connection is still making it difficult to send and receive messages, so this report is a bit briefer than I would otherwise prefer.  Would there be any value in having a technician from Hughes come out to the site to evaluate the link?

In our first few days at the MDRS, Crew 149 has experienced a propane leak, loss of power for nearly 24 hours, water rationing, and interrupted/intermittent comms.  A nominal stateroom has been designated to accommodate the presence of our unofficial eighth crew member, "Murphey."

Aside from that, however, I believe that things are going well.  I am pleased with the matter-of-fact way that my crew has reacted to these challenges.  As one crew member put it, "These are not problems: These are opportunities for us to demonstrate our resilience and resourcefulness!"  Per ardua ad astra, indeed.

Respectfully submitted,

Paul Bakken
Commander, MDRS Crew 149